OpenStudy (anonymous):

I really need help: solve the logarithmic equation ln squareroot of x+5 = 3

OpenStudy (anonymous):

ln square root of x+5 = 3 rewrite: ln(x+5)^(1/2) = 3 square root is also something raised to the half power. which is the same as: (1/2)ln(x+5) = 3 bring the exponent to the front get rid of the (1/2) by multiplying 2 on both sides of the equation. ln(x+5)=6 hope you can figure it out from there if you need an exact numerical answer.